# BiLO: Bilevel Local Operator Learning for PDE Inverse Problems. Part II: Efficient Uncertainty Quantification with Low-Rank Adaptation

### v1

@@ 2.1 Bi-Level Local Operator for Bayesian Inference
Figure 1 panel (C) illustrates the lower-level problem: we train a neural network to approximate the local operator at different $\theta$, e.g. $\theta_{1}$ and $\theta_{2}$. The local operator at $\theta_{1}$ is approximated by a neural network $u(x,\theta_{1};W)$ and is obtained by minimizing the local operator loss $\mathcal{L}_{\text{LO}}(\theta_{1},W)$.
Since we require local operators at different $\theta$ during sampling (e.g., $\theta_{1}$ and $\theta_{2}$), the neural network must be retrained for each $\theta$ to obtain corresponding weights $W(\theta_{1})$ and $W(\theta_{2})$ The weight matrices in the hidden layers of the neural network have dimension $p\times p$, where $p$ is the number of hidden neurons and is usually a large number. Updating the full set of weights $W$ at each sampling step can be computationally expensive, especially when many sampling steps are required. We can reduce the computational cost by using the low-rank adaptation (LoRA) technique [38], which assumes that the change between the weights is low-rank; that is, $W^{*}(\theta_{2})=W^{*}(\theta_{1})+AB$, where $A\in\mathbb{R}^{p\times r}$ and $B\in\mathbb{R}^{r\times p}$ and $r$ is much smaller than $p$. LoRA drastically reduces the number of trainable parameters in the neural network without significantly affecting the performance of the neural network. Further details on the network architecture and the LoRA method are provided in Sections 2.2 and 2.5
@@
Figure 1: A schematic for solving the Bayesian PDE inverse problem using BiLO and LoRA. (A) At the upper level, we sample the PDE parameters $\theta$ from the posterior distribution using gradient-based MCMC. (B) The full operator $u(x,\theta)$ (gray surface) and its local approximation $u(x,\theta_{1})$ and $u(x,\theta_{2})$ (blue surface), which satisfies vanishing residual and residual gradient conditions. The local operators suffices to compute the gradient of the potential energy $U(\theta)$ at $\theta_{1}$ and $\theta_{2}$. (C) At the lower level, we train the neural network to approximate the local operator (blue surface) at different $\theta$. As $\theta$ changes, only low rank adaptation (LoRA) is used to update the neural network weights $W$.
@@ 2.2 Architecture
The representation of the local operator $u(\mathbf{x},\theta;W)$ is based on a multi-layer perceptron (MLP) neural network. We define an embedding layer that maps the inputs $\mathbf{x}\in\Omega\subset\mathbb{R}^{d}$ and the PDE parameters $\theta\in\mathbb{R}^{m}$ to a high-dimensional vector $\mathbf{y}\in\mathbb{R}^{p}$, using an affine transformation followed by an element-wise nonlinear activation function (e.g., tanh) $\sigma$:
$\mathbf{y}=\sigma(\mathbf{W}^{(0)}\mathbf{x}+\mathbf{R}\theta+\mathbf{b}^{(0)}),$ (7)
where $\mathbf{W}^{(0)}\in\mathbb{R}^{p\times d}$ is the embedding matrix applied to $\mathbf{x}$, $\mathbf{R}\in\mathbb{R}^{p\times m}$ is the embedding matrix applied to $\theta$, and $\mathbf{b}^{(0)}\in\mathbb{R}^{p}$ is the bias vector.
We use an MLP that takes $\mathbf{y}$ as input, and define $m(\mathbf{x},\theta;W)$ as its raw output:
$\mathbf{h}^{(0)}$ $=\mathbf{y},$
$\mathbf{h}^{(l)}$ $=\sigma\left(\mathbf{W}^{(l)}\mathbf{h}^{(l-1)}+\mathbf{b}^{(l)}\right),\quad\text{for }l=1,\dots,L,$
$m(\mathbf{x},\theta;W)$ $=\mathbf{W}^{(L+1)}\mathbf{h}^{(L)}+\mathbf{b}^{(L+1)},$
where $\mathbf{W}^{(l)}\in\mathbb{R}^{d_{l}\times d_{l-1}}$, $\mathbf{b}^{(l)}\in\mathbb{R}^{d_{l}}$ are the weights and biases for layer $l$. The output dimension is $d_{L+1}=1$. For simplicity, we assume all hidden layers have the same dimension as the embedding dimension $p$, i.e., $d_{l}=p$ for $l=1,\dots,L$. $W=\{\mathbf{W}^{(l)},\mathbf{b}^{(l)}\}_{l=0}^{L+1}$ denotes the full set of trainable parameters.
A final transformation can be applied to the output of the MLP to enforce the boundary condition or the initial condition [40, 41]. For example, if the PDE is defined on a unit interval $[0,1]$ with Dirichlet boundary conditions $u(0)=u(1)=0$, the BiLO solution can be represented as
$u(\mathbf{x},\theta;W)=m(\mathbf{x},\theta;W)(1-\mathbf{x})\mathbf{x},$ (8)
Alternatively, the boundary condition can be enforced by an additional loss term.
A key design choice in BiLO is that the embedding matrix $\mathbf{R}$ should be non-trainable to prevent it from collapsing to zero. Otherwise, $\mathcal{L}_{\rm rgrad}(W,\theta)$ can be made zero simply by setting $\mathbf{R}$ to be 0. $\mathbf{R}$ can be randomly initialized in the same way as the other weights [42]. We describe the architecture in detail in this section—although standard except for the embedding layer—to facilitate the explanation of Low-Rank Adaptation in Section 2.5.
2.3 Review of Markov Chain Monte Carlo (MCMC) Methods
In this section, we provide a quick overview of the sampling methods that appeared in this work, including the Metropolis-Hastings (MH) algorithm and the Hamiltonian Monte Carlo (HMC). These methods are Markov Chain Monte Carlo (MCMC) methods that generate samples from a target distribution by constructing a Markov chain whose stationary distribution is the target distribution.
We present these methods independent of the BiLO framework. This is assuming that we have the parameter-to-solution map $\theta\mapsto u(\cdot,\theta)$, and we can compute the potential energy $U(\theta)$ and the gradient $\nabla_{\theta}U(\theta)$ accurately. With this assumption, in this section, we temporarily drop the dependence of the potential energy $U$ on the neural network weights $W$. We will reintroduce the dependence on $W$ in Section 2.4 when we describe how to incorporate these sampling methods in the BiLO framework.
Metropolis-Hastings (MH) Algorithm
The Metropolis-Hastings (MH) algorithm [15] is a classical and simple MCMC method. It requires a proposal distribution $Q(\theta^{\prime}|\theta)$ to propose new samples $\theta^{\prime}$ given the current sample $\theta$. The proposal is then accepted or rejected based on the changes in the potential energy. Algorithm 1, given below, summarizes the MH algorithm for sampling from the potential energy $U(\theta)$. For simple proposal distributions, such as Gaussian or uniform distributions, MH is easy to implement and does not require computing the gradient of the potential energy. However, it can be inefficient as the acceptance rate may be low [16]. In this work, the MH algorithm is coupled with an accurate numerical PDE solver, and serves primarily as a reference method to assess the accuracy of alternative sampling approaches.
Algorithm 1 Metropolis-Hastings (MH) Algorithm
1:Initial state $\theta^{(0)}$, proposal distribution $Q(\theta^{\prime}|\theta)$, number of iterations $N$

### v2

@@ 2.1 Method Overview
Figure 1 panel (C) illustrates the lower-level problem and LoRA: we train a neural network to approximate the local operator at different $\theta$, e.g. $\theta_{1}$ and $\theta_{2}$. The local operator at each $\theta_{i}$ is approximated by a neural network $u(x,\theta_{i};W)$ and is obtained by minimizing the local operator loss $\mathcal{L}_{\text{LO}}(\theta_{i},W)$. Instead of updating all the weights $W$ at each $\theta_{i}$, we use LoRA to update the weights for different $\theta_{i}$, which significantly reduces the number of trainable parameters and thus the computational cost.
@@
Figure 1: A schematic for solving the Bayesian PDE inverse problem using BiLO and LoRA. (A) At the upper level, we sample the PDE parameters $\theta$ from the posterior distribution using gradient-based MCMC. (B) The full operator $u(x,\theta)$ (gray surface) and its local approximation $u(x,\theta_{1})$ and $u(x,\theta_{2})$ (blue surface), which satisfies vanishing residual and residual-gradient conditions. The local operator suffices to compute the gradient of the potential energy $U(\theta)$ at $\theta_{1}$ and $\theta_{2}$. (C) At the lower level, we train the neural network to approximate the local operator (blue surface) at different $\theta$. As $\theta$ changes, only low rank adaptation (LoRA) is used to update the neural network weights $W$.
2.2 Example 1: Nonlinear Poisson Equation
In this section, we compare B-BiLO with BPINN [yangBPINNsBayesianPhysicsinformed2021] on a nonlinear Poisson problem to demonstrate the advantages of B-BiLO: physical uncertainty quantification, stable sampling with larger step sizes, and reduced computational cost. We consider the 1D nonlinear Poisson equation [yangBPINNsBayesianPhysicsinformed2021]:
$\lambda u_{xx}(x)+k\tanh(u(x))=f(x),\quad x\in(-0.7,0.7)$ (8)
where $\lambda=0.01$, and both $f$ and the boundary conditions are derived from the ground-truth solution $u^{\rm GT}(x)=\sin^{3}(6x)$ with $k^{\rm GT}=0.7$. The goal is to infer $k$ from observations of $u$, with the prior that $k$ is uniformly distributed on $[0.4,1]$. A key feature of this problem is that the solutions are odd: both $u$ and u(x)-u(-x) solve the PDE, and thus $u(0)=0$ for all $k$ in the range. Therefore, there should be no uncertainty at the origin.
Figure 2: Inference results for the nonlinear Poisson problem. (a) Noisy data, GT solution, mean and std of inferred $u$ using B-BiLO (b) results using BPINN with $\sigma_{f}=0.1$. (c) Posterior distribution of the PDE parameter $k$ using BiLO and BPINN with different $\sigma_{f}$, compared with the reference (Metropolis-Hastings with numerical solver). (d) Accuracy-efficiency comparison: x-axis shows the -log of the std. at the orgin, larger values indicate better physical uncertainty quantification. y-axis shows the effective sample size (ESS) per second of wall time, larger values indicate more efficient sampling. As $\sigma_{f}$ decreases, BPINN becomes more accurate but less efficient. Overall B-BiLO is more accurate and efficient at uncertainty quantification.
Figure 2 compares B-BiLO and BPINN on the nonlinear Poisson problem. Panel (a) shows the noisy data, GT, mean and 2 std. region of the inferred solution $u$ from B-BiLO. The samples of $u$ satisfy the PDE accurately—evidenced by nearly zero posterior standard deviation at the origin. Panel (b) shows BPINN with $\sigma_{f}=0.1$, where the nonzero uncertainty at the origin indicates that individual samples are not solving the PDE accurately. Panel (c) presents the posterior of the PDE parameter $k$. The reference distribution is obtained using Metropolis–Hastings sampling (which is generally less efficient than HMC) combined with a accurate numerical solver. B-BiLO closely matches the reference distribution, while BPINN remains accurate only for small $\sigma_{f}$; larger $\sigma_{f}$ reduces fidelity. Panel (d) compares accuracy and efficiency. The x-axis ($\log_{10}$ of the std at origin) measures physical consistency—larger values indicate more accurate PDE solution. While the y-axis measures the efficiency using the Effective Sample Size (ESS, which quantifies the number of independent samples) per second. For each $\sigma_{f}$, we cross validated the step size $\delta t$ = 0.01, 0.005, 0.001, 0.0005 and report the best ESS/sec. As $\sigma_{f}$ decreases (color from yellow to blue), BPINN becomes more accurate but less efficient, due to the stability constraint on the step size. B-BiLO achieves both high accuracy and high efficiency simultaneously, without tuning $\sigma_{f}$. Additional visualizations are provided in SM.
Figure 3 illustrates the accuracy and efficiency of B-BiLO with LoRA of varying ranks. Panel (a) compares the posterior distribution of $k$ for LoRA ranks 4 and 8 against Full FT, demonstrating that LoRA does not compromise accuracy. Panel (b) reports the relative efficiency of LoRA (rank 8) versus Full FT, plotted as ratios with respect to Full FT. The x-axis shows the number of neurons per layer (depth fixed at 6), indicating increasing network size. The right y-axis shows the relative maximum memory usage, which remains consistently below 1, confirming LoRA’s reduced memory footprint. The left y-axis (blue line) shows the relative average wall time per sample. For small networks, LoRA is slower than Full FT (ratio above 1), while for large networks it becomes faster (ratio below 1). This reflects a trade-off: LoRA reduces the number of trainable parameters but introduces computational overhead and may require more iterations to reach comparable accuracy. As network size grows, the speedup due to reduced trainable parameters outweighs the cost of the overhead and additional iterations.
Figure 3: Accuracy and efficiency of B-BiLO with LoRA of different ranks. (a) Comparison of the posterior distribution of $k$ obtained by B-BiLO with LoRA of different ranks and Full FT. Diamond markers denote the MAP estimates, and bars indicate the 95% highest-density posterior region (HDPR). (b) Relative efficiency of LoRA (rank 8) as ratios with respect to Full FT. The x-axis shows the number of neurons per layer; the left y-axis (blue line) shows the relative average time per sample, and the right y-axis shows the relative maximum memory usage during sampling.
2.3 Example 2: Inferring Patient-Specific Tumor Growth Parameters from MRI Data
We consider a real-world application of B-BiLO for uncertainty quantification of patient-specific parameters GBM growth models using patient MRI data in 2D [zhangPersonalizedPredictionsGlioblastoma2025, balcerakIndividualizingGliomaRadiotherapy2025, ezhovLearnMorphInferNewWay2023, scheufeleFullyAutomaticCalibration2021, lipkovaPersonalizedRadiotherapyDesign2019]. Glioblastoma (GBM) is a highly aggressive brain tumor with poor prognosis, where tumor cells infiltrate beyond lesions visible on MRI, limiting the effectiveness of current treatment strategies and leading to recurrence, thereby underscoring the need for personalized prediction. The infiltration and proliferation of tumor cells can be modeled by reaction–diffusion equations [swansonQuantitativeModelDifferential2000]. Here, we use the formulation presented in [zhangPersonalizedPredictionsGlioblastoma2025]. Let $\Omega$ denote the 2D brain domain derived from MRI data, and the normalized tumor cell density $u(\mathbf{x},t)$ satisfy the Fisher–KPP equation with Neumann boundary conditions:
$\frac{\partial u}{\partial t}=D\bar{D}\gradient\cdot(P(\mathbf{x})\gradient u)+\rho\bar{\rho}u(1-u)\quad\text{in }\Omega$ (9)
where $P$ depends on the tissue distribution (e.g., white and grey matter) obtained from the MRI data, $\bar{D}$ and $\bar{\rho}$ are known patient-specific characteristic parameters based on the data, and $D$ and $\rho$ are the unknown nondimensionalized parameters that we aim to infer.
We consider two regions of interest in the tumor, the whole tumor (WT) region and the tumor core (TC) region. Let $\hat{\mathbf{y}}^{s}$, $s\in\{\rm WT,TC\}$ be indicator functions of the WT and TC regions, which can be obtained from the MRI data and serves as the observational data in the inverse problem. We define our predicted segmentations to be regions in which the tumor cell density $u$ at the nondimensional $t=1$ lies above a certain threshold $u_{c}^{s}$. The likelihood function is defined based on error between the predicted segmentation $\mathbf{y}^{s}$ and the observed segmentation $\hat{\mathbf{y}}^{s}$. We aim to find the posterior distribution of the parameters $(D,\rho,u_{c}^{\rm WT},u_{c}^{\rm TC})$. Details of the setup and validation with synthetic data are provided in the SM (see also [zhangPersonalizedPredictionsGlioblastoma2025]).
Figure 4 summarizes the inference results for one patient case. In panel (a), the posterior distributions of the parameters show that the model parameters are well identified with moderate uncertainty, and the results from B-BiLO with LoRA rank 8 closely match those from Full FT. Panel (b) shows the posterior mean and standard deviation of the tumor cell density $u$, respectively. In panel (c), the MAP segmentation is seen to closely match the observed WT and TC regions on the MRI, and the predicted infiltration margin (where $u=1\%$) extends slightly beyond the visible lesion, suggesting possible microscopic invasion. In panel (d), the standard deviation of the tumor cell density $u$ is shown. Because the data are noisy and the boundary between the TC and WT regions is poorly defined, the estimated uncertainty is higher in this transition zone. Although no direct ground truth exists for the true tumor cell density in patients, the predictive capability of the model can be indirectly validated using recurrence data [zhangPersonalizedPredictionsGlioblastoma2025, balcerakIndividualizingGliomaRadiotherapy2025, balcerakPhysicsRegularizedMultiModalImage2024]. This example illustrates the capability of B-BiLO to handle real patient data for risk-aware treatment planning; applications to full 3D patient datasets are left for future work.
Figure 4: Inference of patient-specific GBM growth parameters using B-BiLO with LoRA rank 8. (a) Posterior distribution of the parameters $D,\rho,u_{c}^{\rm WT},u_{c}^{\rm TC}$ and the prior (dashed), compared with Full FT. (b) Posterior mean of the tumor cell density $u$. (c) Ground truth MRI data with segmented WT and TC regions (filled), the MAP predicted segmentation (contour), and the predicted infiltration margin (where $u=1\%$). (d) Posterior standard deviation of the tumor cell density $u$.
2.4 Example 3: Inferring Stochastic Rates from Particle Data
We consider a problem motivated by inferring the dynamics of gene expression from static images of mRNA molecules in cells [milesInferringStochasticRates2024, milesIncorporatingSpatialDiffusion2025]. The steady-state spatial distribution of these molecules can be modeled by the following boundary value problem with a singular source:
Here, the solution $u(x)$ represents the intensity of a spatial Poisson point process describing the locations of mRNA particles in a simplified 1D domain, and $\delta(x-z)$ is the Dirac delta function representing a point source of mRNA at location $z$. The parameter $\lambda$ is the dimensionless birth rate (transcription) of mRNA at a specific gene site $z$, while $\mu$ is a degradation rate, and the boundary condition describes export across the nuclear boundary.
Given $M$ snapshots of mRNA locations $\{q_{i}^{j}\}$ for $j=1,\ldots,M$, $i=1,\ldots,N_{j}$, where $N_{j}$ is the number of particles in snapshot $j$, the goal is to infer the kinetic parameters $\lambda$ and $\mu$. Since the particle locations follow a Poisson point process, we can construct a likelihood function directly from their positions. The resulting log-likelihood is [milesInferringStochasticRates2024]:
$l(\lambda,\mu;q_{i}^{j})=M\int_{\Omega}u(x;\lambda,\mu)dx-\sum_{j=1}^{M}\sum_{i=1}^{N_{j}}\log u(q_{i}^{j};\lambda,\mu).$
This particular likelihood function, which is standard for spatial Poisson processes, consists of two terms: the sum of the log-intensity at each observed particle’s location and a term penalizing the total expected number of particles in the domain, $\int_{\Omega}u(x)dx$. The solution to the PDE is continuous, but its derivative is discontinuous at the source $z$; we handle this singularity by adapting the cusp-capturing PINN [tsengCuspcapturingPINNElliptic2023]. The details are explained in the SM.
In Fig.5(a), we show the 90% highest-density posterior region (HDPR) and the MAP estimates of $\lambda$ and $\mu$ obtained using the reference method, B-BiLO with LoRA (rank 4), and B-BiLO with Full FT. Both fine-tuning strategies closely match the reference, with relative errors in the MAP estimates and HDPR area below 4% (details in SM). The slight overestimation of $\lambda_{\rm MAP}$ and $\mu_{\rm MAP}$ likely arises from the inexact PDE solution, while the small discrepancies between Full FT and LoRA are attributable to variability in finite MCMC sampling. The inset shows the HDPR and MAP of the solution $u(x)$, which has a cusp at the source location $z=0.5$ and is accurately captured by B-BiLO. In Fig. 5(b), we show the ground truth (GT), prior distributions, and posterior distributions of $\lambda$ and $\mu$.
Figure 5: Inference of gene expression parameters $\lambda$ (birth rate) and $\mu$ (degradation rate) using B-BiLO. (a): The 90% highest-density posterior region (HDPR) and maximum a posteriori (MAP) estimates of the solution obtained using B-BiLO with LoRA rank 4, Full FT, and the reference method (MH with numerical solver). Inset: 90% HDPR and MAP of the solution $u(x)$. (b): Ground truth (GT), prior, and posterior distributions for $\lambda$ and $\mu$, with posterior sampled via B-BiLO using LoRA rank 4.
2.5 Example 4: Darcy Flow Problem
Our framework can be generalized to inferring unknown functions in the PDE, as described in Methods (Sec. 4.5). We consider the following 2D Darcy flow problem on the unit square $\Omega=[0,1]\times[0,1]$ with Dirichlet boundary conditions:
[…]
where $f(\mathbf{x})=10$. The unknown spatially varying diffusion coefficient $D(\mathbf{x})$ is represented by a truncated 64-term Karhunen–Loève expansion with standard normal priors on the coefficients, modeling high- and low-diffusivity regions (3 to 12) across the domain (Details are provided in the SM) [huangEfficientDerivativefreeBayesian2022].
The reference result is obtained by MH sampling with numerical solutions computed on a $\times61$ grid. Noisy observations are collected at $21\times 21$ evenly spaced grid points, and the residual loss is evaluated at $61\times 61$ collocation points. The noise level is set to $\sigma_{d}=0.01$, corresponding to approximately 10% of the maximum value of the solution $u$. Figure 6 illustrates the posterior inference of the spatially varying diffusion coefficient $D(\mathbf{x})$ in the 2D Darcy flow problem. Panel (a) shows the ground truth (GT) diffusion coefficient. Panel (b) presents the point-wise posterior means of $D(\mathbf{x})$ obtained using B-BiLO with LoRA rank 4, which captures the key spatial features of the GT. Panel (c) quantifies the uncertainty in $D(x)$ along a fixed 1D slice of the domain, indicated by the dashed line in (a), showing the region of mean $\pm$ two standard deviations. The relative L2 error of the B-BiLO posterior mean with respect to the reference is 6.2%. Additional visualizations are provided in the SM.
Figure 6: Posterior inference of the spatially varying diffusion coefficient $D(\mathbf{x})$ in the 2D Darcy flow problem. (a) The ground truth $D(\mathbf{x})$; (b) The posterior mean of $D(\mathbf{x})$ obtained using B-BiLO with LoRA rank 4; (c) Mean $\pm$ 2 standard deviations of the posterior $D(\mathbf{x})$ evaluated along a 1D slice of the domain (dashed line in (a) and (b)).
3 Discussion
In this paper, we have introduced a novel framework (B-BiLO) for efficiently solving Bayesian PDE inverse problems. Our approach enforces strong PDE constraints and integrates gradient-based MCMC methods, significantly enhancing computational efficiency and accuracy in parameter inference and uncertainty quantification. Through numerical experiments, we demonstrated that our method provides superior performance compared to Bayesian Physics-Informed Neural Networks (BPINNs). Notably, our approach avoids the computational challenges of directly sampling high-dimensional Bayesian neural network weights and mitigates issues arising from ill-conditioned PDE residual terms. Moreover, by incorporating LoRA, we substantially reduce computational overhead during the fine-tuning phase without compromising the accuracy or robustness of Bayesian inference. Our results on various problems highlight the effectiveness of our method. A promising direction for future work is extending the B-BiLO framework to three-dimensional PDEs and inverse problems with high-dimensional parameter spaces, such as full-field identification problems in geophysics or medical imaging. Tackling these challenges will likely require integration with more efficient sampling algorithms, improved neural network architectures tailored to complex PDEs, and faster fine-tuning strategies to maintain scalability.
4 Methods
@@ 4.1 Architecture
The local operator $u(\mathbf{x},\theta;W)$ is represented by a standard multi-layer perceptron (MLP). The network input is the concatenation of the coordinate and the PDE parameter, $h^{(0)}=[\mathbf{x};\theta]\in\mathbb{R}^{d+m}$. The hidden layers are defined by $h^{(l)}=\sigma(W^{(l)}h^{(l-1)}+b^{(l)})$ for $l=1,\dots,L$, where $W^{(l)}\in\mathbb{R}^{p\times d_{l-1}}$ and $b^{(l)}\in\mathbb{R}^{p}$, with $d_{0}=d+m$ and $d_{l}=p$ for all hidden layers, and $\sigma$ is the activation function ($\tanh$ in this work). The network output is $m(\mathbf{x},\theta;W)=W^{(L+1)}h^{(L)}+b^{(L+1)}$, where $W^{(L+1)}\in\mathbb{R}^{1\times p}$ and $b^{(L+1)}\in\mathbb{R}$. The collection of all trainable parameters is $W=\{W^{(l)},b^{(l)}\}_{l=0}^{L+1}$. A final transformation can be applied to the output of the MLP to enforce the boundary condition or the initial condition [dongMethodRepresentingPeriodic2021, sukumarExactImpositionBoundary2022]. For example, if the PDE is defined on a unit interval $[0,1]$ with Dirichlet boundary conditions $u(0)=u(1)=0$, the BiLO solution can be represented as $u(x,\theta;W)=m(x,\theta;W)(1-x)x$. Alternatively, the boundary condition can be imposed by additional loss terms [raissiPhysicsinformedNeuralNetworks2019]. This MLP architecture is presented as a basic example and to facilitate the explanation of Low-Rank Adaptation in Section 4.4. More advanced architectures can also be incorporated, including residual connections [heDeepResidualLearning2015], random Fourier features [wangEigenvectorBiasFourier2021], and other modified MLP designs [wangUnderstandingMitigatingGradient2021].
4.2 Algorithm
The key component of our method is to compute the gradient of the potential energy $U(\theta)$. In BiLO, the potential energy depends on the weights of the neural network and is denoted as $U(\theta,W)$. At the optimal weights $W^{*}(\theta)$, the local operator $u(\mathbf{x},\theta;W^{*}(\theta))$, the potential energy $U(\theta,W^{*}(\theta))$, and its gradient $\nabla_{\theta}U(\theta,W^{*}(\theta))$ are expected to approximate the exact local operator $u(\mathbf{x},\theta)$, the true potential energy $U(\theta)$, and the true gradient $\nabla_{\theta}U(\theta)$, respectively. Therefore, each time $\theta$ is updated, we must solve the lower level problem. The complete B-BiLO algorithm is detailed in Algorithm 1.
In practice, we solve the lower level problem to some tolerance $\epsilon$ using our subroutine LowerLevelIteration in Algorithm 2. It is written as a simple gradient descent method, but in practice, we can use more advanced optimization methods such as Adam [kingmaAdamMethodStochastic2017]. In the HMC, we call the subroutine LowerLevelIteration each time before computing the gradient of the potential energy $\nabla_{\theta}U(\theta,W)$.
Algorithm 1 B-BiLO